# A Generalized Adaptive Joint Learning Framework for High-Dimensional Time-Varying Models

### v1

@@ Appendix A Proof of Theoretical Results
$S=\{j:\|B_{j}^{\ast}\|_{F}>0\},\qquad S^{c}=\{1,\dots,p\}\setminus S,$
@@
The terms $\mathcal{R}_{g,S}(\Delta)$ and $\mathcal{R}_{f,C}(\Delta)$ involve only the active groups and active changepoints. Since $\mathcal{R}_{g,S}(\Delta)$ involves at most $s=|S|$ groups of size $MK$, and $\mathcal{R}_{f,C}(\Delta)$ involves at most $|C|$ fused differences, Cauchy-Schwarz and the boundedness of the weights imply
@@
Proof of Proposition 2 (Validity of Hierarchical Screening)
Proof.
[…]
1. The KKT Violation Condition
The gradient at the solution $\hat{\theta}$ is:
[…]
2. Bounding Each Term
$\|\text{Term (III)}\|_{F}\leq C_{\text{noise}}\sqrt{\frac{\log p}{N}}.$
3. The Contradiction
Now consider the KKT inequality (30). The RHS (Penalty Threshold) is:
Combining these, we have:
$\text{RHS}=\lambda_{g}\hat{w}_{g,j}\quad(\text{small}).$
$\|\nabla_{\bm{B}_{j}}\mathcal{L}(\hat{\theta})\|_{F}>\lambda_{g}\hat{w}_{g,j}$
@@ 3. Consistency of Variance Estimation
Appendix B Appendix B: Computational Implementation of Slope Fusion (Hierarchical Strategy)
@@
However, simultaneously solving for variable selection (Group Lasso) and effect segmentation (Slope Fused Lasso) for all $p$ covariates in the ultra-high-dimensional regime is computationally burdensome. Therefore, as discussed in Remark 2, we implement this via a hierarchical regularization strategy.

### v2

@@ 3. The Contradiction
$\mathcal{S}=\{j:\|B_{j}^{\ast}\|_{F}>0\},\qquad\mathcal{S}^{c}=\{1,\dots,p\}\setminus\mathcal{S},$
@@
The terms $\mathcal{R}_{g,S}(\Delta)$ and $\mathcal{R}_{f,C}(\Delta)$ involve only the active groups and active changepoints. Since $\mathcal{R}_{g,S}(\Delta)$ involves at most $s=|\mathcal{S}|$ groups of size $MK$, and $\mathcal{R}_{f,C}(\Delta)$ involves at most $|C|$ fused differences, Cauchy-Schwarz and the boundedness of the weights imply
@@ 3. Consistency of Variance Estimation
Appendix B Detailed Justification for the Hierarchical Strategy
@@
However, simultaneously solving for variable selection (Group Lasso) and effect segmentation (Slope Fused Lasso) for all $p$ covariates in the ultra-high-dimensional regime is computationally burdensome. Therefore, as discussed in Section 2.6, we implement this via a hierarchical regularization strategy.